Electric fields can be used to pull electrons out of metals. To remove an electron from tungsten, the electric field must do 4.5 eV of work. Suppose that the distance over which the electric field acts is 3.4 nm. Calculate the minimum force that the field must exert on the electron being removed.

Grade:upto college level

1 Answers

Navjyot Kalra
8 years ago
We have to find out the minimum force (F) that the field must exert on the electron being removed.
To obtain the minimum force (F), substitute 4.5 eV for W and 3.4 nm for s in the equation F = W/s,
F = W/s
= (4.5 eV/3.4 nm)
= (4.5 eV/3.4 nm) (1.6×10-19 J/1 eV) (1 nm/10-9 m)
= 2.1×10 -10 J/m
= 2.1×10 -10 N …… (3)
From equation (3) we observed that, the minimum force (F) that the field must exert on the electron being removed would be 2.1×10 -10 N.